OpenStudy (anonymous):

Solve: 1/b = (6/5b)+1

OpenStudy (lgbasallote):

solve by LCM which is 5b 5 = 6 + 5b 5b = 5-6 5b = -1 b = -1/5

OpenStudy (anonymous):

5/5b = 6/5b +1 -1 = 1/5b 5b = -1/1 b = -1/5